My question is:
find the second coordinates of the points with first coordinates 0 and 1.
k(x) = 5(3x)
Can you help me with this?
Thank you.

Answer by stanbon(75887)

find the second coordinates of the points with first coordinates 0 and 1.
k(x) = 5(3x)
--------------
The points have the form (x,k(x))
----
If x is 0, k(0) = 5(3*0) = 0
giving you a 2nd coordinate of 0.
---------------
If x = 1, k(1) = 5(3*1) = 15
giving you a 2nd coordinate of 15.
======================================
Cheers,
Stan H.